OpenStudy (anonymous):

What is the lim (1-cos theta)/(2sin^2 theta) as theta approahes 0

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0}\frac{1-\cos \theta}{2\sin^2 \theta}=\lim_{\theta \rightarrow 0}\frac{1-\cos \theta}{2(1- \cos^2 \theta)}=\lim_{\theta \rightarrow 0}\frac{1-\cos \theta}{2(1- \cos \theta)(1+\cos \theta)}=\]\[=\frac{1}{4}\]

OpenStudy (anonymous):

Thank you very much